A spaceship takes off vertically from rest with an acceleration of 27.5 m/s2.27.5 m/s2. What magnitude of force FF is exerted on a 54.0 kg54.0 kg astronaut during takeoff? Express FF as a multiple of the astronaut's weight ww on Earth.
Answer #1

a = 27.5 m/s² M = 54 kg N F =ma = (54) (27.5) La = 1485 N. W = mg = (54) (9.81) =529.74 Newton Multiple astronauts weight w